On a horizontal axis whose unit is the meter, a linear load ranging from 0 to 1 ma a linear load distribution = 2 nC / m.determine the modulus of the electric field created by the previous loaded bar at the point A of abscissa 2m (we have to find the relation between l, which is the distance between the elementary bar and the point A and x which sweeps the segment [0: 1]

1
Expert's answer
2021-07-25T09:16:09-0400

"dE=k\dfrac{dq}{x^2}\\
dq=\lambda{dx}\\
\lambda=2nC/m\\
dq=2dxnC\\
dE=k\dfrac{2dx}{x^2}\\
E=2k\int_1^2\dfrac{dx}{x^2}\\
E=2k(\frac{-1}{x})_1^2=k\times10^{-9}N/C\\
E=8.99\times10^9\times10^{-9}N/C\\
E=8.99N/C\\"